OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x - 2y - z, if x = 2, y = 3, and z = -4

OpenStudy (whpalmer4):

Evaluate \[4x-2y-z\]\[x=2,y=3,z=-4\] Plug in the values. What do you get?

OpenStudy (jdoe0001):

\(\bf 4x - 2y - z\qquad x = {\color{red}{ 2}}\quad y = {\color{red}{ 3}}\quad z = {\color{red}{ -4}}\\ \quad \\ 4x - 2y - z\implies 4\cdot {\color{red}{ 2}}-2\cdot {\color{red}{ 3}}-1\cdot {\color{red}{ -4}}\)